# Volumetric medical image segmentation through dual self-distillation in U-shaped networks

### v2

@@ 4 Discussion
The present study investigated the application of self-distillation for medical image segmentation. The goal of this study was to understand whether self-distillation could further improve the segmentation performance of a single U-shaped network to its maximum extent possible by utilizing the knowledge of the deepest encoder and decoder in addition to the ground-truth labels, by using very few additional trainable parameters and without the need of a large pretrained model as in the case of knowledge distillation. Minimizing the KL-divergence in the loss function between the deepest encoder/decoder and shallower encoders/decoders has a regularization effect which helps the model generalize better to unseen data.
@@
The most closely related prior work to DSD is MISSU [36], which applies self-distillation to a transformer-based TransUNet [49], incorporating multi-scale fusion blocks for 3D medical image segmentation. It is worth mentioning that our DSD framework has several advantages over MISSU: (i) DSD is highly adaptable and designed to be compatible with any U-shaped backbone (including VNet, nnU-Net, UNETR, Swin UNETR), while MISSU is tightly integrated with the TransUNet, thereby making it less versatile. (ii) DSD provides a more thorough internal knowledge transfer involving two complementary pathways (encoder and decoder-side distillation), while MISSU relies on a single-path self-distillation process focused on multi-scale feature fusion. (iii) The added learnable bottleneck modules in DSD incur negligible increase in the number of trainable parameters and training time, while the series of atrous convolutions in MISSU significantly increases the number of trainable parameters. (iv) DSD alleviates gradient vanishing issues with deep supervision which helps ensure that shallow layers can benefit directly from both the ground truth and deeper layers, while MISSU does not have a direct supervision at multiple levels that can lead to slower convergence and potentially less effective gradient flow. (v) Quantitative experiments with the MISSU attached to the UNETR framework demonstrate that our DSD framework with UNETR is superior as it obtains a higher Dice score with lower training parameters on both MMWHS and Hippocampus datasets. In summary, while both DSD and MISSU contribute valuable techniques for improving U-shaped networks, DSD’s advantages lie in its flexibility, lower computational burden, and innovative dual self-distillation process.
In our future work, we plan to investigate the effect of shape priors from the ground-truth labels on the DSD framework to further improve our segmentation results.
@@ 5 Conclusion
In this paper, we introduced a novel dual self-distillation framework which could be attached into any U-shaped backbone for volumetric medical image segmentation. We incorporated our DSD framework into UNETR, nnU-Net, VNet and Swin UNETR and evaluated it on various benchmark datasets with promising results. Our results demonstrated that DSD is a efficient and generalized training strategy that could further boost the segmentation performance of these U-shaped networks.

### v3

@@ 4 Discussion
The present study investigated the application of self-distillation for medical image segmentation. The goal of this study was to understand whether self-distillation could further improve the segmentation performance of a single U-shaped network to its maximum extent possible by utilizing the knowledge of the deepest encoder and decoder in addition to the GT Labels, by using very few additional trainable parameters and without the need of a large pretrained model as in the case of knowledge distillation. Minimizing the KL-divergence in the loss function between the deepest encoder/decoder and shallower encoders/decoders has a regularization effect which helps the model generalize better to unseen data.
@@
The observed discrepancies in the performance improvements of our proposed DSD framework across different datasets (MMWHS, MSD-BraTS, and MSD-Hippocampus) and backbone networks are expected due to inherent differences in dataset complexity, modality, and anatomical structures. For instance, the MSD-BraTS multi-modal MRI dataset exhibits high variability in tumor morphology and complexity, making the segmentation inherently more challenging and hence leading to greater potential improvements with DSD attached, compared to simpler, single-modality datasets such as MSD-Hippocampus with T1w MRI and MMWHS CT datasets. These variations illustrate the broad applicability and effectiveness of our DSD framework. Additionally, in Table 4, the observed discrepancies in performance stem from the varying anatomical complexity, size, and image characteristics of each dataset. In MMWHS (CT modality), multiple large cardiac substructures with clearer boundaries create more opportunities for inter- and intra-structure knowledge transfer, leading to pronounced gains with the DSD framework. By contrast, the Hippocampus dataset (T1w MRI) presents smaller, more localized structures with subtler intensity gradients, limiting the available learning signal and reducing the magnitude of improvement. Consequently, while the DSD framework consistently benefits all tested datasets, the degree of enhancement depends on these factors, explaining the stronger results in MMWHS and more modest gains for Hippocampus.
The most closely related prior work to DSD is MISSU [36], which applies self-distillation to a transformer-based TransUNet [49], incorporating multi-scale fusion blocks for 3D medical image segmentation. It is worth mentioning that our DSD framework has several advantages over MISSU: (i) DSD is highly adaptable and designed to be compatible with any U-shaped backbone (including VNet, nnU-Net, UNETR, Swin UNETR), while MISSU is tightly integrated with the TransUNet, thereby making it less versatile. (ii) DSD provides a more thorough internal knowledge transfer involving two complementary pathways (encoder and decoder-side distillation), while MISSU relies on a single-path self-distillation process focused on multi-scale feature fusion. (iii) The added learnable bottleneck modules in DSD incur negligible increase in the number of trainable parameters and training time, while the series of atrous convolutions in MISSU significantly increases the number of trainable parameters. (iv) DSD alleviates gradient vanishing issues with deep supervision which helps ensure that shallow layers can benefit directly from both the ground truth and deeper layers, while MISSU does not have a direct supervision at multiple levels that can lead to slower convergence and potentially less effective gradient flow. (v) Quantitative experiments with the MISSU attached to the UNETR framework demonstrate that our DSD framework with UNETR is superior as it obtains a higher Dice score with a lower number of training parameters on both MMWHS and Hippocampus datasets. In summary, while both DSD and MISSU contribute valuable techniques for improving segmentation performance, DSD’s advantages lie in its flexibility, lower computational burden, and innovative DSD process.
One of the limitation of our approach is its dependency on the U-shaped network structure which means the benefits may not directly translate to other architectures without significant adaptation. DSD also relies on a carefully chosen teacher–student pairing to deliver precise, high-quality guidance and in scenarios where the deepest layers do not provide optimal representations, the improvement might be limited. Additionally, we observed that there are some outlier predictions for some of the complex substructures even after DSD has been attached to the U-shaped backbone. Hence, in our future work, we plan to investigate the effect of shape priors from the GT Labels on the DSD framework to further improve our segmentation results, especially for smaller complex structures.
@@ 5 Conclusion
In this paper, we introduced a novel DSD framework which could be attached into any U-shaped backbone for volumetric medical image segmentation. We incorporated our DSD framework into UNETR, nnU-Net, VNet and Swin UNETR and evaluated it on various benchmark datasets with promising results. Our results demonstrated that DSD is an efficient and general training strategy that could further boost the segmentation performance of these U-shaped networks.